# ORMNet: Object-centric Relationship Modeling for Egocentric Hand-object Segmentation

### v1

@@ Abstract
Egocentric hand-object segmentation (EgoHOS) is a brand-new task aiming at segmenting the hands and interacting objects in the egocentric image. Although significant advancements have been achieved by current methods, establishing an end-to-end model with high accuracy remains an unresolved challenge. Moreover, existing methods lack explicit modeling of the relationships between hands and objects as well as objects and objects, thereby disregarding critical information on hand-object interaction and introducing confusion into algorithms, ultimately leading to a reduction in segmentation performance. To address the limitations of existing methods, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) for EgoHOS. Specifically, based on a single-encoder and multi-decoder framework, we design the Hand-Object Relation (HOR) module to leverage hand-guided attention to capture the correlation between hands and objects and facilitate their representations. Moreover, based on the observed interrelationships between diverse categories of objects, we introduce the Object Relation Decoupling (ORD) strategy. This strategy allows the decoupling of the two-hand object during training, thereby alleviating the ambiguity of the network. Experimental results on three datasets show that the proposed ORMNet has notably exceptional segmentation performance with robust generalization capabilities.
Index Terms:
@@
Egocentric view, hand-object interaction segmentation, attention, relationship modeling.
I Introduction
@@
The advancements in edge computing and computer vision have engendered a growing interest in Head-Mounted Devices (HMD) such as the Apple Vision Pro[1] and Microsoft Hololens 2[2]. Consequently, there is a rising trend in capturing abundant first-person view (FPV) or egocentric images and videos. Compared with data captured from a third-person view (TPV) or an exocentric perspective[3], egocentric data displays the actions and intentions of the HMD wearer more clearly and intuitively, which is more conducive to analyzing human behavior and understanding how humans interact with the environment. Notably, questions concerning the activities of individuals, such as “What is this person doing?” can be more comprehensively addressed through the parsing of egocentric visual data. It is obvious that the hands of HMD wearers and the objects they are interacting with serve as pivotal cues in this analysis.
@@
Therefore, to delve deeply into human behavior encoded within egocentric data at a fine-grained level, Zhang et al.[4] first proposed the Egocentric Hand-Object Segmentation (EgoHOS) task, which aims to perform pixel-level segmentation of the hand and the objects that interact with the hand in egocentric images. This novel segmentation task enables a precise response to the query, “Where are the left and right hands, and what objects are they interacting with?” which can function as an essential algorithm serving as the foundation for various subsequent applications, such as augmented reality/virtual reality (AR/VR)[5, 6], medical assistant[7, 8], robotics[9], embodied AI[10, 11], and metaverse[12].
Figure 1: Structural differences between the previous (a) parallel method, (b) sequential method, and (c) our approach. Compared with previous parallel and sequential methods, the proposed ORMNet can achieve higher segmentation performance with higher efficiency by integrating relationship modeling between “hands and objects (H-O)” as well as “objects and objects (O-O).” The $N,N_{h},$ and $N_{o}$ mean the number of overall categories, the number of hand categories, and the number of object categories, and $N=N_{h}+N_{o}$.
The previous approaches for EgoHOS can be categorized into parallel and sequential methods. Specifically, parallel methods[4, 13, 14, 15, 16] are straightforward, which employ an end-to-end encoder-decoder framework to extract features and predict the mask of all categories simultaneously, as shown in Fig. 1(a). In contrast to parallel methods, sequential approaches[4, 17] employ a step-by-step prediction scheme for the hand and interacting objects, i.e., the input of the latter step to predict the objects’ masks necessitates the previous step’s output for predicting the hand, as depicted in Fig. 1(b).
Although existing methods have demonstrated a satisfied EgoHOS capability, they are accompanied by many drawbacks, as shown in the table in Fig. 1. First, similar to conventional encoder-decoder semantic segmentation networks[18, 19], the intuitive parallel approaches[4] regard hands and objects as independent categories and predict segmentation results for them in an end-to-end manner. While this type of technique demonstrates high efficiency, it fails to capture the relationship between “hands and objects” or among different objects, consequently leading to performance deficiencies. Second, in contrast, the sequential methods[4] specially built for the EgoHOS task demonstrate enhanced performance by incorporating the relationships between hands and interacting objects. However, step-wise training and inference elongate the overall time consumption, thereby reducing the method’s efficiency and constraining its feasibility for real-time applications. Third, the prediction of diverse interacting objects is required in the EgoHOS task, i.e., left-hand objects (objects interact with the left hand), right-hand objects (objects interact with the right hand), and two-hand objects (objects interact with both hands at the same time). This paper observes the coupling relationship between these three categories of objects, i.e., the two-hand object can be regarded as both a left-hand object and a right-hand object. Therefore, it is easy to predict a two-hand object as the left-hand object when the left interaction is more obvious than the right interaction, and the same applies to the right-hand object. Nevertheless, prevailing methods overlook the ambiguity of such relationships, which introduces confusion during network training.
@@
To improve efficiency and effectiveness simultaneously, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) to model the relationship between “hands and objects” as well as “objects and objects,” as shown in Fig. 1(c). On the one hand, we propose the Hand-Object Relation (HOR) module to establish the correlation between the features of the hand and interacting objects. The core underlying concept involves leveraging hand features as prior knowledge and employing hand-guided attention to extract more representative object features. On the other hand, we introduce a new Object Relation Decoupling (ORD) strategy to model the relationship among diverse object categories. Drawing from the insights gained from analyzing two-hand objects, this strategy restricts the prediction to left- and right-hand objects exclusively during the training phase. Then, two-hand objects are generated through a simple yet efficient intersection operation during inference. This strategy can reduce the confusion of the algorithm about the object categories during learning, enabling the network to concentrate more precisely on the interacting objects. The contributions of this paper are as follows:
@@
  • This paper introduces a novel end-to-end ORMNet for the EgoHOS task that achieves state-of-the-art segmentation performance while being highly efficient.
@@
  • To take full advantage of the prominence of hand features, the proposed HOR module models the relationship between hands and objects, helping the network extract more comprehensive hand and object features.
  • Based on the relation of various object categories, the simple yet powerful ORD strategy is proposed to decouple the two-hand objects during training, reducing the ambiguity of the network and further separately highlighting the objects contacted with left or right hands.
@@
  • The experimental results on two datasets demonstrate that the proposed ORMNet achieves exceptional performance and manifests robust generalization capabilities.
II Related Works
II-A Image Segmentation
Image segmentation is a basic task of computer vision, which aims to divide the image into several non-overlapping regions based on certain principles, such as semantic information. Semantic segmentation[20, 21, 22], instance segmentation[23, 24, 25], and panoptic segmentation[26, 27, 28] are three commonly studied segmentation sub-tasks. Image segmentation methods can be divided into traditional methods, convolution neural network (CNN)-based methods, and transformer-based methods in chronological order. Although traditional methods can perform image segmentation with notable interpretability, further refinement is needed to enhance the quality of the segmentation outcomes. The fully connected neural network (FCN)[19] proposed by Jonathan et al. is the cornerstone of CNN segmentation methods. After that, there have been numerous advancements[29] in CNN-based image segmentation methods, such as U-Net[18], Deeplab-series[30, 31], SegNet[32], PSPNet[33], etc.
The development of transformer-based methods for image segmentation is a relatively recent but rapidly evolving field. Dosovitskiy et al. introduced the Vision Transformer (ViT)[34] in 2020, proving that the transformer could be applied directly to images by treating the image as a sequence of patches. Following the success of ViT, several transformer-based methods[35] specially designed for image segmentation have been proposed. Zheng et al. developed the SETR[36] to adopt ViT as a backbone to extract features to eliminate the reliance on FCN, achieving impressive performance. Similarly, Strudel et al. proposed a pure end-to-end Transformer approach - Segmenter[37], which utilized the ViT encoder for capturing global context and a simple point-wise linear decoder to generate precise segmentation masks. Meanwhile, Xie et al. presented the Segformer[13] to integrate hierarchical Transformer encoder with lightweight multi-layer perceptron (MLP) decoders, avoiding complex designs in segmentation methods. Unlike previous work, Cheng et al. proposed the MaskFormer[38], which abandoned the per-pixel classification formulation and instead predicted a series of binary masks associated with class labels. Following that, Cheng et al. introduced the Mask2Former[39], which incorporated masked attention to perform universal (i.e., instance, semantic, and panoptic) segmentation, achieving notable performance. Then, Jain et al. proposed OneFormer[14] to perform universal segmentation tasks with transformers and task-guided queries, which achieved state-of-the-art performance in various segmentation benchmarks. Recently, Kirillov et al. proposed a new impressive work called Segment Anything (SAM)[40]. They built a foundation model for interactive segmentation by training with over 1 billion masks on 11M images.
Our paper utilizes the Swin transformer[41] to extract the features of the input egocentric image and to up-sample the features to the input size[15]. Although using more recent methods may improve performance, this is not the main focus of this paper.
II-B Egocentric Hand-object Segmentation
Zhang et al.[4] recently introduced a novel task known as Egocentric Hand-object Segmentation (EgoHOS), which combines image segmentation and Human-Object Interaction detection (HOI-det). This task redefines the focus of HOI-det from the entire human body to specific, fine-grained parts, i.e., human hands. Its objective is to segment the left hand, right hand, and the objects that interact with the left and right hands in input RGB egocentric images. Concurrently, Zhang et al. [4] introduced the innovative concept of the contact boundary (CB), representing the region where the hands and the objects engage in interaction. This concept explicitly promotes the network focus on the interaction between hands and objects. Furthermore, based on the proposed task and concept, the authors also proposed different types of methods in their work, including parallel methods and sequential methods.
Parallel method. A more direct and naive method is to use an encoder and a decoder to segment and predict hands and objects and use different indexes to imply different categories[4, 13, 14, 15, 16], as shown in Fig. 1(a). Although this parallel method completes the prediction of this task simply and directly, its segmentation accuracy is relatively low and still needs to be improved.
@@ II-B Egocentric Hand-object Segmentation
Sequential method. Since parallel methods lacked targeted design in learning hand-object interaction, Zhang et al. proposed a sequential method[4], as shown in Fig. 1(b). Taking the egocentric image as input, this step-by-step sequential method uses an encoder-decoder first to segment the left and right hands. After that, the segmented left- and right-hand masks are concatenated to the input image to be used to perform the mask prediction of contact boundary and interacting objects. Although the sequential method achieved higher segmentation performance than the parallel method, it lacks efficiency due to the step-by-step process.
@@
In addition, both parallel and sequential methods inadequately explore the relationship between hands and objects and ignore the relationship between objects and objects. Unlike the previous parallel and sequential methods, the proposed ORMNet can achieve end-to-end training and inference with high accuracy due to the relationship modeling between “hands and objects” as well as “objects and objects.”
II-C Attention Mechanism in deep learning
Models using the attention mechanism have exploded in recent years, confirming the effectiveness of the attention mechanism for NLP[42, 43] and computer vision[44, 45] tasks. The primary objective of attention in vision is to mimic the selective function of the human visual cognitive system towards input information, i.e., concentrating on the primary and crucial features of the input visual data. The majority of attention modules utilize distinct methodologies to generate an attention map for feature weighting, consequently adjusting the significance of individual components within the features.
In the early stage, most attention modules are integrated into the CNN-based models. For example, Hu et al. proposed the SENet[46] equipped with channel-wise attention, which used global average pooling to generate the attention weight among channel axial. After that, Woo et al. proposed CAM [47] to improve channel attention by aggregating the descriptors from max-pooling and average-pooling. There are many other methods intended to improve channel-wise attention, e.g., SOCA[48], HA[49], etc. Unlike channel attention, spatial attention is proposed to generate weights for spatial patches of feature maps. For example, Woo et al. proposed SAM [47] to use a convolutional layer to obtain the attention weights.
The development of self-attention and multi-head self-attention (MHSA) mechanisms in transformer-based models has also been widely used in various methods in recent years. Some researchers also make an effort to extend self-attention to cross-attention[50, 51], which allows the model to capture relationships between different modalities of data[52, 53]. The proposed ORMNet utilizes hand-guided cross-attention to model the relationship between the features of hands and objects, allowing more comprehensive feature extraction.
Figure 2: Overall diagram of the proposed ORMNet, which consists of an encoder, bottleneck, multiple decoder branches, Hand-Object Relation (HOR) module, and Object Relation Decoupling (ORD) strategy. The algorithm is utilized to predict the hands, right-hand objects (RO), left-hand objects (LO), and two-hand objects (TO). The ORD strategy processes the initial (Ini.) ground truth (GT) to generate new supervision during data preparation, and it also processes the predictions (Pred.) of diverse categories of interacting objects during inference.
III Methodology
@@
This section introduces the proposed method in detail from top to bottom, including task description (in Sec. III-A), overall architecture (in Sec. III-B), Object Relation Decoupling Strategy (in Sec. III-E), Hand-Object Relation module (in Sec. III-D), and training and inference (in Sec. III-F).
III-A Task Description
@@
The EgoHOS dataset [4] includes the label of the hands, $1^{st}$-order interacting objects, and $2^{nd}$-order interacting objects. In this paper, we focus on the segmentation of hands and $1^{st}$-order interacting objects, i.e., objects that are directly contacted by hands. Given an egocentric image I, the goal of the EgoHOS task is to predict the masks of the following entities: left hand $\textbf{M}_{H}^{L}$, right hand $\textbf{M}_{H}^{R}$, left-hand objects $\textbf{M}_{O}^{L}$ (objects directly interacting with the left hand), right-hand objects $\textbf{M}_{O}^{R}$ (objects directly interacting with the right hand), and two-hand objects $\textbf{M}_{O}^{T}$ (objects interacting with both hands simultaneously).
III-B Overall Architecture

### v2

@@ Abstract
Egocentric hand-object segmentation (EgoHOS) is a promising new task aiming at segmenting hands and interacting objects in egocentric images. Although EgoHOS has the potential to enable various applications, current methods struggle to achieve both high performance and end-to-end optimization simultaneously. Moreover, existing approaches fail to fully leverage hand cues to assist the interacting-object segmentation and overlook the coupled relationships between diverse interacting-object categories, resulting in performance deficiencies. To address these limitations, this paper proposes a novel Object-centric Relationship Modeling Network (ORMNet) to fulfill end-to-end and effective EgoHOS by modeling relationships between hands and objects as well as objects and objects. Specifically, a Hand-Object Relation (HOR) module is introduced to capture the correlation between hands and objects, which uses hand features to guide the network to extract more distinguishing interacting-object features. Besides, we find the coupling relations between diverse interacting-object categories and design the Object Relation Decoupling (ORD) strategy to disentangle them, emphasizing learning of the interaction between hands and objects and reducing the confusion of interacting-object classification. In-domain experiments show that ORMNet has notably exceptional segmentation performance compared with state-of-the-art methods, while out-of-domain experiments further exhibit its robust generalization capability. The project is available at https://github.com/yuggiehk/ORMNet/ .
keywords:
@@
Egocentric perception , hand-object segmentation , cross-attention , relationship modeling
journal: Pattern Recognition
[…]
[label1]organization=Department of Electrical and Electronic Engineering, The Hong Kong Polytechnic University, addressline=Hung Hum, city=Kowloon, country=Hong Kong SAR
1 Introduction
@@
The advancements in edge computing and computer vision [1, 2, 3] have engendered a growing interest in Head-Mounted Devices (HMD) such as the Apple Vision Pro [4] and Microsoft Hololens 2 [5]. Consequently, there is a rising trend in capturing abundant first-person view (FPV) or egocentric images and videos [6, 7]. Compared with data captured from a third-person view (TPV) or an exocentric perspective [8], egocentric data displays more detailed, intuitive, and clear actions and intentions of the HMD wearer, which is more conducive to analyzing human behavior and understanding how people interact with the environment. Notably, questions concerning the activities of individuals, such as “What is this person doing?” can be more comprehensively addressed in detail through the parsing of egocentric visual data. It is obvious that the hands of HMD wearers and the objects they are interacting with serve as pivotal cues in this analysis.
@@
Therefore, to delve deeply into human behavior encoded within egocentric data at a fine-grained level, Zhang et al. [9] first proposed a new task, i.e., Egocentric Hand-Object Segmentation (EgoHOS), which aims to perform pixel-level segmentation of hands and objects that are interacting with hands in egocentric images. This novel segmentation task enables a precise response to the query, “Where are the left and right hands, and what objects are they interacting with?” which can function as an essential algorithm serving as the foundation for various subsequent applications, such as augmented reality/virtual reality (AR/VR) [10, 11], medical assistive systems [12, 13], and embodied AI [14, 15].
Figure 1: (a) parallel methods [9, 16, 17, 18, 19, 40], (b) sequential methods [9], and (c) our approach. Comparisons of segmentation performance (d) on the EgoHOS in-domain and out-of-domain test sets show that our ORMNet can achieve significant segmentation and generalization performance with end-to-end optimization by integrating the HOR module and ORD strategy.
The previous approaches for EgoHOS can be categorized into parallel and sequential methods. Specifically, some unified segmentation methods [9, 16, 17, 18, 19] can be directly used as parallel methods in this new task, which adopt an end-to-end framework to first extract image features and then simultaneously predict all categories of masks, as shown in Fig. 1 (a). However, due to the newly-emerging EgoHOS task, applying such generic segmentation methodologies directly to this task lacks targeted design of relation modeling, resulting in suboptimal effectiveness. In contrast to parallel methods, sequential approaches [9] are specially designed to form a step-by-step prediction scheme for hands and interacting objects, i.e., the input of the latter step to predict the objects’ masks necessitates the previous step’s output for predicting hands, as depicted in Fig. 1 (b). Although the effectiveness of sequential methods surpasses that of parallel methods, they are unable to achieve end-to-end optimization, which reduces the efficiency of the network. Therefore, achieving both high performance and end-to-end optimization in this new EgoHOS task remains an unresolved issue.
Furthermore, regardless of variations in interactions and environments, it is easier to recognize hands in egocentric images by identifying the shape and texture [20, 7]. Conversely, the segmentation of interacting objects presents significant challenges due to the reliance on discerning interaction with hands rather than visual features. Therefore, the dominant role of hands in the EgoHOS task underscores the benefit of incorporating hands to assist in predicting interacting objects. However, as mentioned above, parallel methods [9, 21, 22] treat hands and interacting objects as completely independent categories for prediction, while sequential methods [9] use predicted hand masks as input for interacting-object segmentation. Both approaches fail to subtly and explicitly model the dominance of hands in predicting interacting objects, thereby leading to performance deficiencies.
In addition, the EgoHOS task requires the prediction of three types of interacting objects determined by the hand(s) they interact with, namely, left-hand objects, right-hand objects, and two-hand objects, as described in detail in Sec. 3.1. We observe that the two-hand object encompasses attributes of both the left- or right-hand object because they are interacting with both left- and right-hand simultaneously. However, most existing parallel and sequential approaches treat different types of interacting objects as separate, which requires the network to learn “whether the object is interacting with both hands at the same time,” introducing additional confusion into the training process and resulting in inaccurate results.
@@
To address the shortcomings of the existing methods above, this paper proposes a novel Object-centric Relationship Modeling Network (ORMNet) to model the relationships between “hands and objects” as well as “objects and objects”, which can achieve end-to-end optimization with significant performance, as shown in Fig. 1 (c). On the one hand, we propose the Hand-Object Relation (HOR) module to establish the correlation between the features of the hand and interacting objects. The core underlying concept involves leveraging hand features as prior knowledge and employing hand-guided attention to extract more representative interacting-object features. On the other hand, based on the observations about interacting objects, we introduce a novel Object Relation Decoupling (ORD) strategy to disentangle the interrelationships among diverse interacting-object categories. This strategy restricts the prediction of the left- and right-hand objects exclusively during training, and another type of object (two-hand objects) is generated through an ingenious and pure mathematical intersection operation during inference. This strategy can help the algorithm emphasize the interaction and reduce the confusion of the algorithm about the interacting-object classification during learning. Experimental results on in-domain and out-of-domain test sets demonstrate the state-of-the-art performance of ORMNet along with robust generalization capabilities, as shown in Fig. 1 (d).
The contributions of this paper are as follows:
  • 1.
@@
    This paper introduces a novel ORMNet for the EgoHOS task, achieving state-of-the-art segmentation performance while fulfilling end-to-end optimization.
  • 2.
@@
    To take full advantage of the prominence of hand features, ORMNet incorporates the HOR module to model the relationship between hands and interacting objects, helping the network extract more representative object features.
  • 3.
    Based on the observed relation between various interacting-object categories, an innovative ORD strategy is proposed to decouple them, decreasing the concentration on classification and emphasizing interaction learning of the algorithm.
  • 4.
@@
    The thorough experimental results demonstrate that the proposed ORMNet achieves exceptional performance and manifests robust generalization capabilities.
2 Related Work
This section mainly introduces the related work of this paper. Since the EgoHOS task is strongly related to Egocentric Hand-Object Interaction (EgoHOI), we will first introduce EgoHOI (Sec. 2.0.1) and then refine to EgoHOS (Sec. 2.0.2).
2.0.1 Egocentric Hand-Object Interaction
Egocentric images and videos offer a clear window into how humans physically engage with their surrounding environment and the objects using their hands. In recent years, the interpretation and analysis of egocentric visual data have earned increasing attention from the research community, especially in the field of Egocentric hand-object interaction (EgoHOI) [23, 24, 25] due to the release of various large-scale datasets [26, 24, 27, 28, 29].
Aiming to solve the EgoHOI task, certain efforts [30] used additional multi-modality information, such as gaze [31] and text [32, 33] cues, to enhance the learnable representations. Furthermore, self-supervised diagrams [34] were developed to effectively leverage the underlying visual information. Although egocentric visual data is inherently far more limited in quantity compared to the abundant third-person perspective (TPV) data, some works [35, 36] proposed to learn the joint representations and transfer the view-agnostic knowledge from TPV to FPV. However, the data requirements are strict, time-consuming, and labor-intensive, which needs videos capturing the same action from different perspectives. In addition, EgoHOI can be utilized as a basic algorithm to perform other tasks, such as action recognition [20, 37, 38] and action anticipation [39]. While many transformer-based approaches have made breakthrough progress on the EgoHOI task, the sole prediction of bounding boxes around the interacting objects lacks the granularity of fine-grained analysis.
2.0.2 Egocentric Hand-Object Segmentation
Different from EgoHOI, EgoHOS is a novel problem proposed by Zhang et al. [9] to perform fine-grained pixel-level segmentation for hands and interacting objects. The final prediction output encompasses a set of categories, including the left hand, right hand, left-hand objects, right-hand objects, and two-hand objects. Furthermore, the authors also proposed different types of methods in their work, including parallel methods and sequential methods.
Parallel methods. Some general segmentation methods [9, 16, 17, 18, 19, 40] can be directly used as parallel methods to perform EgoHOS task. These methods take egocentric images as input, and directly obtain predictions of all categories after feature extraction and mask generation, as shown in Fig. 1 (a). Although these methods can complete the EgoHOS task in an end-to-end manner, they are not effective due to the lack of task-specific design.
@@ 2.0.2 Egocentric Hand-Object Segmentation
Sequential methods. Zhang et al. proposed a series of sequential methods [9] tailored for the EgoHOS task, as shown in Fig. 1 (b). Taking the egocentric image as input, these step-by-step sequential methods first segment the left and right hands. After that, the segmented left- and right-hand masks are concatenated to the input image to perform the mask prediction of interacting objects. Although sequential methods achieve higher segmentation performance than the parallel method, the step-by-step process is unable to achieve end-to-end optimization, which limits the applications in the real world.
@@
In addition, both parallel and sequential methods inadequately explore the relationship between hands and interacting objects, which does not emphasize the leading role of the hand [20] in the EgoHOS task, thereby degrading the segmentation performance. Finally, parallel and sequential methods regard the different categories of interacting objects as independent and ignore the potential relationships between them, causing unnecessary confusion in classifying the interacting-object categories. Unlike the previous parallel and sequential methods, the proposed ORMNet can achieve end-to-end optimization with high accuracy due to the relationship modeling between “hands and objects” as well as “objects and objects,” as shown in Fig. 1 (c).
3 Methodology
@@
This section introduces the proposed ORMNet in detail from top to bottom, including task description (in Sec. 3.1), overall architecture (in Sec. 3.2), HOR module (in Sec. 3.3), ORD strategy (in Sec. 3.4), and training and inference (in Sec. 3.5).
3.1 Task Description
@@
This section first delineates the detailed description of the newly-emerging EgoHOS task. The EgoHOS task includes the segmentation of left- and right-hands, $1^{st}$-order interacting objects, and $2^{nd}$-order interacting objects. Following the work proposed by Zhang et al. [9], we focus on the segmentation of hands and $1^{st}$-order interacting objects, i.e., objects that are directly contacted by hands. Given an egocentric image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict the masks of the following entities: hands $\textbf{M}_{H}$, left-hand objects $\textbf{M}_{O}^{L}$ (objects directly interacting with the left hand), right-hand objects $\textbf{M}_{O}^{R}$ (objects directly interacting with the right hand), and two-hand objects $\textbf{M}_{O}^{T}$ (objects interacting with both hands simultaneously).
Figure 2: Overall diagram of the proposed end-to-end ORMNet, which incorporates the HOR module and ORD strategy to effectively predict hands, left-hand objects (LO), right-hand objects (RO), and two-hand objects (TO). The ORD strategy consists of data preparation (left), training (middle), and inference (right).
3.2 Overall Architecture